# Beyond Ensemble Averages: Leveraging Climate Model Ensembles for Subseasonal Forecasting

### v3

@@ D.1 U-Net details
Figure 24: U-Net architecture with input channels = $C$. $C$ is the number of input channels, which in our case equals the number of ensemble members plus all climate data.
Our U-Net differs from the original U-Net by modifying the first 2D convolutional layer after input. Since our input channels can be different when we choose a different subset of features or different ensemble (NCEP-CFSv2 or NASA-GMAO), this 2D convolutional layer is used to transform our input with $C$ channels into a latent representation with 64 channels. The number of channels $C$ depends on which ensemble we are using, and what task we are performing. For example, for precipitation tasks using the NCEP-CFSv2 ensemble, the input channels include $24$ ensemble members, $5$ lagged observations, $4$ other observational variables, $8$ principal components of SSTs, and 24 positional encoding, resulting in $65$ channels in total. For temperature tasks using the NASA-GMAO ensemble, there are only $11$ ensemble members, and we don’t include SSTs information, hence there are only $44$ channels in total. The other following layers use the same configurations with the standard U-Net [49].
We also perform careful hyperparameter tuning for the U-NET. In particular, we run 10-fold cross validation on our training set, and use grid search for tuning learning rate, batch size, number of epochs, and weight decay. Since we use different loss functions for different forecast tasks and different number of input channels for NCEP-CFSv2 and GMAO-GMAO ensemble, we run hyperparameter tuning with the same cross validation scheme separately for these tasks. For instance, for precipitation regression, we choose from 100, 120, 150, 170, 200, 250 epochs; batch size may be equal to 8, 16, 32; learning rate values are chosen from 0.0001, 0.001, 0.01; weight decay can be 0, 1e-3, 1e-4. In case of NCEP-CFSv2 precipitation regression, the optimal parameters are 170 epochs, batch size 16, learning rate 0.0001, and weight decay 1e-4. For temperature regression using the same data, the best parameters are: 100 epochs, batch size 16, learning rate 0.001, and weight decay 1e-3. For tercile classification of precipitation, the best parameters are 80 epochs (we chose from 60, 70, 80, 90, 100 epochs during classification), batch size 8, learning rate 0.001, and weight decay 1e-4.
@@ D.2 Random Forest Quantile Regressor details
Figure 25: Illustration of a Random Forest, which serves as a visual aid for the ensuing discussion on existing and proposed methods. Quantile Regression Forests outputs the empirical $\alpha$ quantiles of the collection of all responses in all the leaves associated with $\mathbf{x}_{t,l}$ (marked with a star for each tree).
@@ D.3 Stacking model details
The stacking model is a simple one layer neural network with 100 hidden neurons and sigmoid activation function for regression and softmax for classification. We use an implementation from Scikit-learn library [48]. We choose 100 neurons based on the stacking model performance on the validation data (we also try 50, 75, 100 and 120 neurons). The stacking model demonstrates stable performance in general, but with 100 neurons it usually achieves the best results. We use ”lbfgs” optimizer from quasi-Newton methods for the regression tasks, and SGD optimizer for classification tasks.
@@ Appendix E Preprocessing Details
There are more than 100,000 SSTs locations available. We extract the top eight principal components. Principal component analysis fits on train part and then is applied for the rest of the data. In the case of U-Net, we deal with PCs of SSTs by adding additional input channels that are constant across space with each channel corresponding to one of PCs. Random forest can use PCs from SSTs directly with no special preprocessing.
@@
We apply channel-wise min-max normalization to the input features at each location based on training part of the dataset in the case of U-Net. As for normalization of the true values, min-max normalization is applied for precipitation and standardization is applied for temperature. This choice affects the final layer of U-Net model too: for the precipitation regression task, the sigmoid activation is used, and no activation function is applied for temperature regression. For the stacking model, we apply min-max normalization to both input and target values.

### v4

@@ D.1 U-Net details
Figure 24: U-Net architecture with input channels = $C$. $C$ is the number of input channels, which, in our case, equals the number of ensemble members plus all climate data.
Our U-Net differs from the original U-Net by modifying the first 2D convolutional layer after input. Since our input channels can be different when we choose a different subset of features or different ensemble (NCEP-CFSv2 or NASA-GMAO), this 2D convolutional layer is used to transform our input with $C$ channels into a latent representation with 64 channels. The number of channels $C$ depends on which ensemble we are using and what task we are performing. For example, for precipitation tasks using the NCEP-CFSv2 ensemble, the input channels include $24$ ensemble members, $5$ lagged observations, $4$ other observational variables, $8$ principal components of SSTs, and 24 positional encodings, resulting in $65$ channels in total. For temperature tasks using the NASA-GMAO ensemble, there are only $11$ ensemble members, and we don’t include SSTs information, hence there are only $44$ channels in total. The other following layers use the same configurations with the standard U-Net [49].
We also perform careful hyperparameter tuning for the U-NET. In particular, we run a 10-fold cross-validation on our training set, and use grid search for tuning learning rate, batch size, number of epochs, and weight decay. Since we use different loss functions for different forecast tasks and different numbers of input channels for NCEP-CFSv2 and GMAO-GMAO ensemble, we run hyperparameter tuning with the same cross-validation scheme separately for these tasks. For instance, for precipitation regression, we choose from 100, 120, 150, 170, 200, and 250 epochs; batch size may be equal to 8, 16, 32; learning rate values are chosen from 0.0001, 0.001, 0.01; weight decay can be 0, 1e-3, 1e-4. In case of NCEP-CFSv2 precipitation regression, the optimal parameters are 170 epochs, batch size 16, learning rate 0.0001, and weight decay 1e-4. For temperature regression using the same data, the best parameters are 100 epochs, batch size 16, learning rate 0.001, and weight decay 1e-3. For tercile classification of precipitation, the best parameters are 80 epochs (we chose from 60, 70, 80, 90, and 100 epochs during classification), batch size 8, learning rate 0.001, and weight decay 1e-4.
@@ D.2 Random Forest Quantile Regressor details
Figure 25: Illustration of a random forest, which serves as a visual aid for our discussion. Quantile regression forests outputs the empirical $\alpha$ quantiles of the collection of all responses in all the leaves associated with $\mathbf{x}_{t,l}$ (marked with a star for each tree).
@@ D.3 Stacking model details
The stacking model is a simple one-layer neural network with 100 hidden neurons and a sigmoid activation function for regression and softmax for classification. We use an implementation from Scikit-learn library [48]. We choose 100 neurons based on the stacking model performance on the validation data (we also try 50, 75, 100 and 120 neurons). The stacking model demonstrates stable performance in general, but with 100 neurons it usually achieves the best results. We use the “lbfgs” optimizer from quasi-Newton methods for the regression tasks, and the SGD optimizer for classification tasks.
@@ Appendix E Preprocessing Details
There are more than 100,000 SSTs locations available. We extract the top eight principal components. Principal component analysis fits on the train part and then is applied to the rest of the data. In the case of U-Net, we deal with PCs of SSTs by adding additional input channels that are constant across space, with each channel corresponding to one of PCs. Random forest can use PCs from SSTs directly with no special preprocessing.
@@
We apply channel-wise min-max normalization to the input features at each location based on the training part of the dataset in the case of U-Net. As for normalization of the true values, min-max normalization is applied for precipitation, and standardization is applied for temperature. This choice affects the final layer of the U-Net model, too: for the precipitation regression task, a sigmoid activation is used, and no activation function is applied for temperature regression. For the stacking model, we apply min-max normalization to both input and target values.